# A localisation phase transition for the catalytic branching random walk

### v1

@@ 4.2 Checking Assumption (A1)
$\kappa Q_{x}=\Big{(}\frac{2\lambda_{x}}{1+\lambda_{x}}-1\Big{)}(1+\lambda_{x})\delta_{x}+\frac{1}{1+\lambda_{x}}\mathbb{E}[(1+\lambda_{x+\Delta})\delta_{x+\Delta}]=(\lambda_{x}-1)\delta_{x}+\frac{1}{1+\lambda_{x}}\mathbb{E}[(1+\lambda_{x+\Delta})\delta_{x+\Delta}].$ (4.5)
@@
$\kappa Q_{0}(\mathbb{Z}^{d})=\lambda_{0}-1+\frac{\gamma+\lambda}{1+\lambda_{0}}=\lambda_{0}-\frac{\lambda_{0}-\lambda}{1+\lambda_{0}}.$
@@
$\lambda_{0}-\frac{\lambda_{0}-\lambda}{1+\lambda_{0}}=\lambda+(\lambda_{0}-\lambda)\cdot\frac{\lambda_{0}}{1+\lambda_{0}}>\lambda-(\lambda_{0}-\lambda)\cdot\frac{2d-1+2d\lambda}{2d(1+\lambda)}\geq\lambda-\frac{\lambda_{0}-\lambda}{2d(1+\lambda)},$
@@
because $2d-1+2d\lambda\geq 2d-1\geq 1$. Thus the maximum of $x\mapsto\kappa Q_{x}(\mathbb{Z}^{d})$ is reached at $x=0$ and because, by definition, $\kappa=Q_{0}(E)$ (see (4.1); this is why we chose this value for $\kappa$), we get
$\max_{x\in\mathbb{Z}^{d}}Q_{x}(E)=1.$ (4.7)
This concludes the proof of (4.7). From (4.6), we also get that
$\min_{x\in\mathbb{Z}^{d}}Q_{x}(\mathbb{Z}^{d})=\frac{\lambda}{\kappa}<1.$ (4.8)
Thus, to prove that (A1) holds with $c=\lambda>0$ and $\beta=2$, it is enough to show that
$\sup_{x\in\mathbb{Z}^{d}}\,\mathbb{E}\big{|}Q_{x}^{\scriptscriptstyle(1)}(\mathbb{Z}^{d})-Q_{x}(\mathbb{Z}^{d})\big{|}^{2}<+\infty,$
@@
which holds because, for all $x\in\mathbb{Z}^{d}$, $Q_{x}^{\scriptscriptstyle(1)}(\mathbb{Z}^{d})$ is almost surely bounded (by $-(1+\lambda_{0})/\kappa$ from below, and by $2(1+\lambda_{0})/\kappa$ from above).
@@ 4.3 Checking Assumption (A’2)
(i) For all $N\geq 0$, $\{x\in\mathbb{Z}^{d}\colon\|x\|\leq N\}$ is finite and thus compact.
@@
$\kappa Q_{x}V\leq\lambda V(x)+1+\max_{\|x\|_{1}\leq 1}Q_{x}V.$ (4.9)
Similarly, for all $\alpha\in(\nicefrac{{1}}{{2}},1)$, for all $x\in\mathbb{Z}^{d}$,
@@
As $\|x\|_{1}\uparrow\infty$, $(\|x\|_{1}+1)^{\alpha}\sim\|x\|_{1}^{\alpha}$. We fix $\varepsilon=\varepsilon(\alpha)>0$ small enough so that $\lambda+\varepsilon<\kappa$ (this is possible because $\lambda<\kappa$ see (4.8)). Then, we choose $L(\varepsilon)\geq 1$ large enough so that $\|x\|_{1}\geq L(\varepsilon)$ implies $(\|x\|_{1}+1)^{\alpha}\leq(1+\varepsilon)\|x\|_{1}^{\alpha}$. With these choices, for all $x\in\mathbb{Z}^{d}$ such that $\|x\|_{1}\geq L(\varepsilon)$,
@@
By Equations (4.9) and (4.10), we get that (ii) holds for any $q=\nicefrac{{1}}{{\alpha}}\in(1,2)$: indeed, one can take $\theta=(\lambda+\varepsilon(\alpha))/\kappa<1$ and $K=\max_{\|x\|_{1}\leq L(\varepsilon(\alpha))}V(x)^{\alpha}$.
@@
$|\kappa Q^{\scriptscriptstyle(1)}_{x}f|=|(2B_{x}-1)(1+\lambda_{x})f(x)+(1-B_{x})(1+\lambda_{x+\Delta})f(x+\Delta)|\leq 2(1+\lambda_{0})\|f\|_{\infty}\leq 2(1+\lambda_{0}).$
@@
$|R_{x}f|\leq\mathbb{E}[|R^{\scriptscriptstyle(1)}_{x}f|]\leq 2/\kappa\quad\text{ and }\quad|Q_{x}f|\leq\mathbb{E}[|Q^{\scriptscriptstyle(1)}_{x}f|]\leq 2(1+\lambda_{0})/\kappa.$
@@
$|Q_{x}f|^{2}\leq\bigg{(}\frac{2(1+\lambda_{0})}{\kappa}\bigg{)}^{2}\leq\frac{4(1+\lambda_{0})^{2}}{\kappa^{2}}\leq\frac{4(1+\lambda_{0})^{2}}{\kappa^{2}}\cdot V(x),$
because $V(x)\geq 1$ for all $x\in\mathbb{Z}^{d}$. Now, for all $r>1$, using first the triangular inequality and then the fact that $x\mapsto x^{r}$ is convex and thus $(a+b)^{r}\leq 2^{r-1}(a^{r}+b^{r})$ for all $a,b\in\mathbb{R}$, we get
$\mathbb{E}\big{[}\big{|}{R}^{\scriptscriptstyle(1)}_{x}\cdot f-R_{x}\cdot f\big{|}^{r}\big{]}\leq\mathbb{E}\big{[}\big{(}\big{|}{R}^{\scriptscriptstyle(1)}_{x}\cdot f\big{|}+\big{|}R_{x}\cdot f\big{|}\big{)}^{r}\big{]}\leq 2^{r-1}\big{(}\mathbb{E}\big{[}\big{|}{R}^{\scriptscriptstyle(1)}_{x}\cdot f\big{|}^{r}\big{]}+\big{|}R_{x}\cdot f\big{|}^{r}\big{)}\leq 2^{r}(2/\kappa)^{r}\leq(\nicefrac{{4}}{{\kappa}})^{r}.$
@@
$\mathbb{E}\big{[}\big{|}{R}^{\scriptscriptstyle(1)}_{x}\cdot f-R_{x}\cdot f\big{|}^{2}\big{]}\leq(1\vee\nicefrac{{4}}{{\kappa}})^{2}\leq(1\vee\nicefrac{{4}}{{\kappa}})^{2}V(x).$
@@
$\leq(1+\lambda_{0})^{q}(V(x)^{\alpha}+V(x+\Delta)^{\nicefrac{{1}}{{q}}})^{q}$
@@
because $q\geq 1$ and thus $x\mapsto x^{q}$ is convex, implying that $(a+b)^{q}\leq 2^{q-1}(a^{q}+b^{q})$. We have proved before that, for all $x\in\mathbb{Z}^{d}$, $V(x+\Delta)\leq V(x)+1$ almost surely. This implies that, almost surely,
@@
$|Q_{x}V^{\nicefrac{{1}}{{q}}}|^{q}\leq\mathbb{E}[|Q^{\scriptscriptstyle(1)}_{x}V^{\nicefrac{{1}}{{q}}}|^{q}]\leq 2^{q+1}(1+\lambda_{0})^{q}V(x)/\kappa.$ (4.12)

### v2

@@ 5.2 Checking Assumption (A’1)
$\kappa Q_{x}$ $=\Big{(}\frac{2\lambda_{x}}{1+\lambda_{x}}-1\Big{)}(1+\lambda_{x})\delta_{x}+\frac{1}{1+\lambda_{x}}\mathbb{E}[(1+\lambda_{x+\Delta})\delta_{x+\Delta}]$
$=(\lambda_{x}-1)\delta_{x}+\frac{1}{1+\lambda_{x}}\mathbb{E}[(1+\lambda_{x+\Delta})\delta_{x+\Delta}].$ (5.4)
@@
$\kappa Q_{0}(\mathbb{Z}^{d})=\lambda_{0}-1+\frac{1+\lambda}{1+\lambda_{0}}=\lambda_{0}-\frac{\lambda_{0}-\lambda}{1+\lambda_{0}}.$
@@
$\lambda_{0}-\frac{\lambda_{0}-\lambda}{1+\lambda_{0}}=\lambda+(\lambda_{0}-\lambda)\cdot\frac{\lambda_{0}}{1+\lambda_{0}}>\lambda+(\lambda_{0}-\lambda)\cdot\frac{2d-1+2d\lambda}{2d(1+\lambda)}\geq\lambda+\frac{\lambda_{0}-\lambda}{2d(1+\lambda)},$
@@
because $2d-1+2d\lambda\geq 2d-1\geq 1$. Thus the maximum of $x\mapsto\kappa Q_{x}(\mathbb{Z}^{d})$ is reached at $x=0$ and because, by definition, $\kappa=Q_{0}(\mathbb{Z}^{d})$ (see (5.1); this is why we chose this value for $\kappa$), we get
$\max_{x\in\mathbb{Z}^{d}}Q_{x}(\mathbb{Z}^{d})=1,$ (5.6)
as needed. Now note that, by definition,
$m_{n+1}P(\mathbb{Z}^{d})=m_{n}P(\mathbb{Z}^{d})+Q_{Y_{n+1}}^{\scriptscriptstyle(n+1)}(\mathbb{Z}^{d}),$
where, for all $x\in\mathbb{Z}^{d}$,
$\mathbb{P}(Y_{n+1}=x|m_{n})=\frac{m_{n}P(x)}{m_{n}P(\mathbb{Z}^{d})}.$
This implies
$m_{n}P(\mathbb{Z}^{d})$ $=(1+\lambda_{0})+\sum_{i=1}^{n}Q_{\xi(i)}^{\scriptscriptstyle(i)}(\mathbb{Z}^{d})$
$=(1+\lambda_{0})+\sum_{i=1}^{n}\mathbb{E}_{i-1}[Q_{\xi(i)}^{\scriptscriptstyle(i)}(\mathbb{Z}^{d})]+\sum_{i=1}^{n}\big{(}Q_{\xi(i)}^{\scriptscriptstyle(i)}(\mathbb{Z}^{d})-\mathbb{E}_{i-1}[Q^{\scriptscriptstyle(i)}_{\xi(i)}(\mathbb{Z}^{d})]\big{)},$ (5.7)
where $\mathbb{E}_{i-1}$ denotes the conditional expectation with respect to $m_{1},\ldots,m_{i-1}$, for all $i\geq 1$. We treat the two sums above separately. First note that, for all $i\geq 1$, by independence of $Q^{\scriptscriptstyle(i)}$ and $m_{1},\ldots,m_{i-1}$,
$\mathbb{E}_{i-1}[\kappa Q^{\scriptscriptstyle(i)}_{\xi(i)}(\mathbb{Z}^{d})]=\sum_{x\in\mathbb{Z}^{d}}\kappa Q_{x}(\mathbb{Z}^{d})\cdot\frac{m_{n}P(x)}{m_{n}P(\mathbb{Z}^{d})}\geq\lambda+(\kappa-\lambda)\frac{m_{n}P(0)}{m_{n}P(\mathbb{Z}^{d})}.$ (5.8)
Indeed, $\kappa Q_{x}(\mathbb{Z}^{d})\geq\lambda$ for all $x\in\mathbb{Z}^{d}$ and $\kappa Q_{0}(\mathbb{Z}^{d})=\kappa$. By definition, $m_{n}P(0)=(1+\lambda_{0})\Pi_{\tau_{n}}(0)$ and $m_{n}P(\mathbb{Z}^{d})=(1+\lambda)N_{\tau_{n}}+(\lambda_{0}-\lambda)\Pi_{\tau_{n}}(0)\leq(1+\lambda_{0})N_{\tau_{n}}$. Thus,
$\liminf_{n\to\infty}\frac{m_{n}P(0)}{m_{n}P(\mathbb{Z}^{d})}\geq\liminf_{n\to\infty}\frac{\Pi_{\tau_{n}}(0)}{N_{\tau_{n}}}\geq\frac{\lambda_{0}-\lambda-1}{\lambda_{0}-\lambda},$
almost surely, by Theorem 1.4 (i) (because $\tau_{n}\to+\infty$ almost surely as $n\to+\infty$). Thus, almost surely,
$\liminf_{n\to\infty}\mathbb{E}_{i-1}[\kappa Q^{\scriptscriptstyle(i)}_{\xi(i)}(\mathbb{Z}^{d})]\geq\lambda+(\kappa-\lambda)\frac{\lambda_{0}-\lambda-1}{\lambda_{0}-\lambda},$
which implies
$\liminf_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{i-1}[\kappa Q_{\xi(i)}^{\scriptscriptstyle(i)}(\mathbb{Z}^{d})]\geq\lambda+(\kappa-\lambda)\frac{\lambda_{0}-\lambda-1}{\lambda_{0}-\lambda}.$ (5.9)
We now control the second sum in (5.7): note that $(\Phi_{n}:=\sum_{i=1}^{n}(Q_{\xi(i)}^{\scriptscriptstyle(i)}(\mathbb{Z}^{d})-\mathbb{E}_{i-1}[Q^{\scriptscriptstyle(i)}_{\xi(i)}(\mathbb{Z}^{d})]))_{n\geq 0}$ is a martingale. Also,
$\sup_{x\in\mathbb{Z}^{d}}\,\mathbb{E}\big{|}Q_{x}^{\scriptscriptstyle(1)}(\mathbb{Z}^{d})-Q_{x}(\mathbb{Z}^{d})\big{|}^{2}<+\infty.$
@@
Indeed, for all $x\in\mathbb{Z}^{d}$, $Q_{x}^{\scriptscriptstyle(1)}(\mathbb{Z}^{d})$ is almost surely bounded (by $(1+\lambda_{0})/\kappa$ from below, and by $2(1+\lambda_{0})/\kappa$ from above). This implies that
$\langle\Phi\rangle_{n}=\sum_{i=1}^{n}\mathbb{E}_{i-1}\big{[}(Q_{\xi(i)}^{\scriptscriptstyle(i)}(\mathbb{Z}^{d})-\mathbb{E}_{i-1}[Q^{\scriptscriptstyle(i)}_{\xi(i)}(\mathbb{Z}^{d})])^{2}\big{]}\leq n\sup_{x\in\mathbb{Z}^{d}}\mathbb{E}\big{|}Q_{x}^{\scriptscriptstyle(1)}(\mathbb{Z}^{d})-Q_{x}(\mathbb{Z}^{d})\big{|}^{2}=\mathcal{O}(n),$
almost surely as $n\to+\infty$. Thus, by the law of large numbers for martingales (see, e.g. [Duf97, Theorem 1.3.15]), $\lim_{n\to\infty}\Phi_{n}/n=0$ almost surely. Using this and (5.9) in (5.7) gives
$\liminf_{n\to\infty}\frac{m_{n}P(\mathbb{Z}^{d})}{n}\geq\frac{\lambda}{\kappa}+\bigg{(}1-\frac{\lambda}{\kappa}\bigg{)}\frac{\lambda_{0}-\lambda-1}{\lambda_{0}-\lambda}.$ (5.10)
For $\liminf_{n\uparrow\infty}\frac{1}{n}\sum_{i=1}^{n}Q_{\xi(i)}(\mathbb{Z}^{d})$, we reason similarly: first note that, for all $n\geq 1$,
$\frac{1}{n}\sum_{i=1}^{n}Q_{\xi(i)}(\mathbb{Z}^{d})=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{i-1}[Q_{\xi(i)}(\mathbb{Z}^{d})]+\frac{1}{n}\sum_{i=1}^{n}\big{(}Q_{\xi(i)}(\mathbb{Z}^{d})-\mathbb{E}_{i-1}[Q_{\xi(i)}(\mathbb{Z}^{d})]\big{)}.$ (5.11)
We look at these two sums separately: for the first sum, note that
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{i-1}[\kappa Q_{\xi(i)}(\mathbb{Z}^{d})]=\frac{1}{n}\sum_{i=1}^{n}\sum_{x\in\mathbb{Z}^{d}}\kappa Q_{x}(\mathbb{Z}^{d})\frac{m_{n}P(x)}{m_{n}P(\mathbb{Z}^{d})}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{i-1}[\kappa Q^{\scriptscriptstyle(i)}_{\xi(i)}(\mathbb{Z}^{d})],$
by (5.8). Thus, by (5.9), we get
$\liminf_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{i-1}[\kappa Q_{\xi(i)}(\mathbb{Z}^{d})]\geq\lambda+(\kappa-\lambda)\frac{\lambda_{0}-\lambda-1}{\lambda_{0}-\lambda}.$ (5.12)
For the second sum in (5.11), we note that $(\hat{\Phi}_{n}:=\sum_{i=1}^{n}(Q_{\xi(i)}(\mathbb{Z}^{d})-\mathbb{E}_{i-1}[Q_{\xi(i)}(\mathbb{Z}^{d})]))_{n\geq 0}$ is a martingale such that
$\langle\hat{\Phi}\rangle_{n}=\sum_{i=1}^{n}\mathbb{E}_{i-1}\big{[}\big{(}Q_{\xi(i)}(\mathbb{Z}^{d})-\mathbb{E}_{i-1}[Q_{\xi(i)}(\mathbb{Z}^{d})]\big{)}^{2}\big{]}\leq 4n\sup_{x\in\mathbb{Z}^{d}}Q_{x}(\mathbb{Z}^{d})^{2}=\mathcal{O}(n).$
Thus, almost surely as $n\uparrow\infty$, $\hat{\Phi}_{n}=o(n)$. Thus, by (5.11) and (5.12), we get that, almost surely as $n\uparrow\infty$,
$\liminf_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}Q_{\xi(i)}(\mathbb{Z}^{d})\geq\frac{\lambda}{\kappa}+\bigg{(}1-\frac{\lambda}{\kappa}\bigg{)}\frac{\lambda_{0}-\lambda-1}{\lambda_{0}-\lambda}.$
Together with (5.10), this implies that (A’1) holds with
$c=\frac{\lambda}{\kappa}+\bigg{(}1-\frac{\lambda}{\kappa}\bigg{)}\frac{\lambda_{0}-\lambda-1}{\lambda_{0}-\lambda}.$ (5.13)
Remark.
Because $\min_{x\in\mathbb{Z}^{d}}Q_{x}(\mathbb{Z}^{d})=\lambda$, it would have been much easier to prove that [MV20, Assumption (A1)] holds with $c=\lambda/\kappa$, but this constant is not large enough for (A’2) to hold. This is the reason why we need to first prove the weak localisation result in Theorem 1.4 (i) before using MVPPs to get a strong localisation result.
@@ 5.3 Checking Assumption (A’2)
(i) For all $N\geq 0$, $\{x\in\mathbb{Z}^{d}\colon\|x\|_{1}\leq N\}$ is finite and thus compact.
@@
$\kappa Q_{x}V\leq\lambda V(x)+1+\max_{\|x\|_{1}\leq 1}\kappa Q_{x}V.$ (5.14)
Similarly, for all $\alpha\in(\nicefrac{{1}}{{2}},1)$, for all $x\in\mathbb{Z}^{d}$ such that $\|x\|_{1}>1$,
@@
As $\|x\|_{1}\uparrow\infty$, $(\|x\|_{1}+1)^{\alpha}\sim\|x\|_{1}^{\alpha}$. We fix $\varepsilon>0$ small enough so that $(\lambda+\varepsilon)/\kappa<c$ (this is possible because $\lambda/\kappa<c$ - see (5.13)). Then, we choose $L(\varepsilon,\alpha)>1$ large enough so that $\|x\|_{1}\geq L(\varepsilon,\alpha)$ implies $(\|x\|_{1}+1)^{\alpha}\leq(1+\varepsilon)\|x\|_{1}^{\alpha}$. With these choices, for all $x\in\mathbb{Z}^{d}$ such that $\|x\|_{1}\geq L(\varepsilon,\alpha)$,
@@
By Equations (5.14) and (5.15), we get that (ii) holds for any $q=\nicefrac{{1}}{{\alpha}}\in(1,2)$: indeed, one can take $\theta=(\lambda+\varepsilon)/\kappa<c$ and $K=\max_{\|x\|_{1}\leq L(\varepsilon,\alpha)}V(x)^{\alpha}/\kappa$.
@@
$|\kappa Q^{\scriptscriptstyle(1)}_{x}f|$ $=|(2B_{x}-1)(1+\lambda_{x})f(x)+(1-B_{x})(1+\lambda_{x+\Delta})f(x+\Delta)|$
$\leq 2(1+\lambda_{0})\|f\|_{\infty}\leq 2(1+\lambda_{0}).$
@@
$|R_{x}\cdot f|\leq\mathbb{E}[|R^{\scriptscriptstyle(1)}_{x}\cdot f|]\leq 2/\kappa\quad\text{ and }\quad|Q_{x}\cdot f|\leq\mathbb{E}[|Q^{\scriptscriptstyle(1)}_{x}\cdot f|]\leq 2(1+\lambda_{0})/\kappa.$
@@
$|Q_{x}\cdot f|^{2}\leq\bigg{(}\frac{2(1+\lambda_{0})}{\kappa}\bigg{)}^{2}\leq\frac{4(1+\lambda_{0})^{2}}{\kappa^{2}}\leq\frac{4(1+\lambda_{0})^{2}}{\kappa^{2}}\cdot V(x),$
because $V(x)\geq 1$ for all $x\in\mathbb{Z}^{d}$. Now, for all $r>1$, using first the triangular inequality and then the fact that $x\mapsto x^{r}$ is convex on $0,\infty)$ and thus $(a+b)^{r}\leq 2^{r-1}(a^{r}+b^{r})$ for all $a,b\in[0,\infty)$, we get
$\mathbb{E}\big{[}\big{|}{R}^{\scriptscriptstyle(1)}_{x}\cdot f-R_{x}\cdot f\big{|}^{r}\big{]}$ $\leq\mathbb{E}\big{[}\big{(}\big{|}{R}^{\scriptscriptstyle(1)}_{x}\cdot f\big{|}+\big{|}R_{x}\cdot f\big{|}\big{)}^{r}\big{]}\leq 2^{r-1}\big{(}\mathbb{E}\big{[}\big{|}{R}^{\scriptscriptstyle(1)}_{x}\cdot f\big{|}^{r}\big{]}+\big{|}R_{x}\cdot f\big{|}^{r}\big{)}$
$\leq 2^{r}(2/\kappa)^{r}\leq(\nicefrac{{4}}{{\kappa}})^{r}.$
@@
$\mathbb{E}\big{[}\big{|}{R}^{\scriptscriptstyle(1)}_{x}\cdot f-R_{x}\cdot f\big{|}^{r}\big{]}\leq(1\vee\nicefrac{{4}}{{\kappa}})^{r}\leq(1\vee\nicefrac{{4}}{{\kappa}})^{2}V(x).$
@@
$\leq(1+\lambda_{0})^{q}(V(x)^{\nicefrac{{1}}{{q}}}+V(x+\Delta)^{\nicefrac{{1}}{{q}}})^{q}$
@@
because $q\geq 1$ and thus $x\mapsto x^{q}$ is convex on $0,\infty)$, implying that $(a+b)^{q}\leq 2^{q-1}(a^{q}+b^{q})$ for all $a,b\in[0,\infty)$. We have proved before that, for all $x\in\mathbb{Z}^{d}$, $V(x+\Delta)\leq V(x)+1$ almost surely. This implies that, almost surely,
@@
$|Q_{x}V^{\nicefrac{{1}}{{q}}}|^{q}\leq\mathbb{E}[|Q^{\scriptscriptstyle(1)}_{x}V^{\nicefrac{{1}}{{q}}}|^{q}]\leq 2^{q+1}(1+\lambda_{0})^{q}V(x)/\kappa^{q}.$ (5.17)